17. Find the 20th term from the last term of the A.P. 3, 8, 13, …, 253.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given A.P. is3, 8, 13, …, 253 Common difference, d= 5. Therefore, we can write the given AP in reverse order as; 253, 248, 243, …, 13, 8, 5 Now for the new AP, first term, a = 253 and common difference, d = 248 − 253 = −5 n = 20 Therefore, using nth term formula, we get, a20 = a+(20−1)d a20 = 253+(19)(−5) a20 = 253−95 a = 158 Therefore, 20th term from the last term of the AP 3, 8, 13, …, 253.is 158.